# IgCraft: A versatile sequence generation framework for antibody discovery and engineering

### v1

@@ 1 Introduction
IgCraft is a generative model that uses a Bayesian Flow Network (BFN) (Graves et al., 2023; Atkinson et al., 2024) to sample paired human antibody sequences under a variety of contexts relevant to therapeutic antibody development. Similar to diffusion models, BFNs perform a denoising process jointly across all sequence positions, decoupling the number of generation steps from sequence length and enforcing no particular generation order amongst the tokens (unlike autoregressive models). Crucially, the BFN’s lack of a specific generation order enables flexible sequence inpainting (conditional generation of subsequences) from a single model trained only to generate full-length sequences, as was demonstrated in ProtBFN (Atkinson et al., 2024) with a sequential monte carlo approach. The ability to perform inpainting for arbitrary sequence regions is especially attractive given the multitude of design scenarios encountered in therapeutic antibody development, including developability optimisation, humanisation, and affinity maturation. As a model that can efficiently perform both unconditional and conditional sequence sampling with the flexibility to incorporate structural data (when available), IgCraft offers substantial practical advantages as a single tool that can be used for a variety of tasks in real-world antibody engineering workflows.
@@ 2.1 Bayesian flow networks
Where is a predicted vector of probabilities over token categories and is an accuracy schedule, playing a similar role to the variance schedule in a diffusion model. In practice, is a neural network that is given the current logits for all sequence positions as input and uses learned relationships between these noisy variables to make a prediction for each token’s value . We use for all experiments as in the original BFN paper (Graves et al., 2023). We also introduce a temperature parameter that scales the network’s output logits before they are converted into probabilities via softmax and use for unconditional sampling and for all other (conditional) sampling tasks. To solve the SDE in (1), we implement a second-order solver similar to that proposed by Xue et al. (2024) and perform sampling in 20 steps. For conditional sampling, we use the particle filtering method outlined in ProtBFN (Atkinson et al., 2024) with 32 particles. More details on IgCraft’s sampling methodology can be found in Appendix A.
@@ 2.3 Datasets and training regime
Training IgCraft consists of three stages. First, each chain’s transformer stack (Figure 1, blue) is pre-trained on unpaired sequences. Then, the model is fine-tuned on paired sequences, for which the pre-trained weights for both transformer stacks are loaded into the network and the interaction blocks (Figure 1, green) are randomly initialized. However, we initialize the bias term in the output gate of each interaction block to a value of , negating its contribution to the token embeddings at the start of fine-tuning and effectively initializing the model as two unpaired sequence models that do not communicate. All weights are updated during this stage of fine-tuning, including the weights initialized from pre-training. Finally, the model is fine-tuned using paired antibody structures as conditioning information for the model’s sequence predictions, initializing the output gate of the structure encoder to . During structure-fine-tuning, the framework regions of input structures are masked stochastically in 50% of training examples to train the model to perform CDR-conditional framework generation. In this stage, only the weights in the structure encoder are updated to ensure that the main trunk retains its capabilities as a pure sequence generative model.
@@ 3.3 Inverse folding and structure-guided sequence design
To assess the model’s ability to conditionally generate antibody sequences in the presence of structural information, we sampled a single sequence for each of the 98 test structures using IgCraft with a temperature of 0.1, providing the antibody and target backbone structures as input to the model’s structure encoder but exposing no sequence information. For comparison, we also performed inverse folding (also using temperature 0.1) using ProteinMPNN (Dauparas et al., 2022), AbMPNN (Dreyer et al., 2023), and Antifold (Høie et al., 2024). While the target was cropped to 128 residues when sampling with IgCraft, the full target chains were provided as input to all other inverse folding methods. To measure performance, we calculate the amino acid recovery for each variable domain region and estimate two properties related to the developability potential of the generated sequences: humanness and solubility. Humanness scores for the generated sequences are obtained using AbNatiV (Ramon et al., 2024) while solubility is scored with CamSol (Sormanni et al., 2015; Rosace et al., 2023), both of which have been supported by experimental validation. Results are shown in Table 3.
@@ 3.4 CDR grafting and humanisation
Scaffolding (”grafting”) of CDRs from mouse or other non-human antibodies into human antibody framework regions is a key task in antibody engineering (Jones et al., 1986; Kim & Hong, 2012), since non-human antibodies are often easier to obtain but can induce pathological anti-drug immune responses when administered as therapeutics (Khazaeli et al., 1994). While traditional CDR grafting workflows typically rely on sequence homology searches to identify human frameworks with high levels of sequence similarity to the input non-human antibody (Kim & Hong, 2012), HuDiff was recently proposed as a ML-driven solution for this task (Ma et al., 2024). HuDiff uses a discrete diffusion model to conditionally generate framework sequences given a set of input CDR sequences, and is fine-tuned specifically on mouse antibodies to sample mutations that improve the AbNatiV humanness score (Ramon et al., 2024) of generated sequences. The authors include impressive experimental evidence demonstrating that the binding properties of an existing high-affinity mouse antibody (<1nM kD) were largely preserved after humanisation. However, HuDiff provides no mechanism for integrating structural information for the input CDRs even though CDR grafting (and motif scaffolding more broadly) is fundamentally a structural task.
@@
Given IgCraft’s strong performance in generating human framework sequences and its capacity to condition on structural information, we sought to evaluate the model’s ability to generate human framework sequences for scaffolding mouse CDRs. We performed conditional sampling with IgCraft using 27 mouse antibodies from SAbDab as input and generating framework sequences for both chains, providing the CDR structures as conditioning information as well as the CDR sequences 2 residues on each side. For each set of mouse CDRs, we also sampled a single paired sequence from HuDiff (Ma et al., 2024) as a benchmark. The effectiveness of CDR grafting is measured in two ways: first, the extent to which the humanness of the antibody is improved after grafting (compared to the parental antibody), and second, the extent to which the binding properties of the antibody are preserved. We evaluated humanness using AbNatiV (Ramon et al., 2024), a machine learning-based approach for humanness scoring, and OASis (Prihoda et al., 2022), an approach that estimates humanness using on 9-mer peptide frequences in OAS. To test how well both grafting methods maintained the structure and binding properties of the wild-type CDRs, we applied AlphaFold3 (Abramson et al., 2024) to fold each generated sequence (and its target, if applicable) with 10 seeds, as well as the 27 original mouse sequences for comparison. For each structure prediction we measure the RMSD of the H-cdr3 loop (superimposing only on the framework region), and for bound antibodies, determine whether the antibody was docked to the correct epitope on the target protein using the widely-applied DockQ score threshold of 0.23 (Mirabello & Wallner, 2024; Abramson et al., 2024). Results are shown in Table 4. We also include an ablation study in Appendix A in which CDR grafting was performed using IgCraft without structural information, demonstrating that providing CDR structures as input significantly improves the model’s ability to propose framework sequences that preserve the structural features of the input CDRs.
@@ A.1.3 Conditional sampling
For conditional sampling tasks with IgCraft we implement the sequential monte carlo (SMC) framework from ProtBFN (Trippe et al., 2022; Atkinson et al., 2024). In general, SMC methods attempt to approximate a posterior distribution over the data given some conditioning information :
In sequential importance resampling, at each step of the sampling trajectory, proposal samples are first drawn from the unconditional model and then re-sampled with replacement under some appropriately normalized likelihood function . The trained BFN - via the SDE in (1) - provides access to the unconditional distribution over token logits . The problem of sequence inpainting considers the task of conditioning generation on a subset of sequence positions with corresponding tokens . Denoting samples of the proposal distribution as , ProtBFN (Atkinson et al., 2024) uses the following form for the conditional likelihood:
@@ A.1.4 Gated residual connections
A key element of IgCraft’s network architecture is the use of gated residual connections at key positions between layers that are updated at different stages of the training process (Figure 1). A standard sigmoid-gated linear unit (GLU) (Dauphin et al., 2016) is defined as follows:

### v2

@@ 1 Introduction
IgCraft is a generative model that uses a Bayesian Flow Network (BFN) (Graves et al., 2023; Atkinson et al., 2025) to sample paired human antibody sequences under a variety of contexts relevant to therapeutic antibody development. Similar to diffusion models, BFNs perform a denoising process jointly across all sequence positions, decoupling the number of generation steps from sequence length and enforcing no particular generation order amongst the tokens (unlike autoregressive models). Crucially, the BFN’s lack of a specific generation order enables flexible sequence inpainting (conditional generation of subsequences) from a single model trained only to generate full-length sequences, as was demonstrated in ProtBFN (Atkinson et al., 2025) with a sequential monte carlo approach. The ability to perform inpainting for arbitrary sequence regions is especially attractive given the multitude of design scenarios encountered in therapeutic antibody development, including developability optimisation, humanisation, and affinity maturation. As a model that can efficiently perform both unconditional and conditional sequence sampling with the flexibility to incorporate structural data (when available), IgCraft offers substantial practical advantages as a single tool that can be used for a variety of tasks in real-world antibody engineering workflows.
@@ 2.1 Bayesian flow networks
Where is a predicted vector of probabilities over token categories and is an accuracy schedule, playing a similar role to the variance schedule in a diffusion model. In practice, is a neural network that is given the current logits for all sequence positions as input and uses learned relationships between these noisy variables to make a prediction for each token’s value . We use for all experiments as in the original BFN paper (Graves et al., 2023). We also introduce a temperature parameter that scales the network’s output logits before they are converted into probabilities via softmax and use for unconditional sampling and for all other (conditional) sampling tasks. To solve the SDE in (1), we implement a second-order solver similar to that proposed by Xue et al. (2024) and perform sampling in 20 steps. For conditional sampling, we use the particle filtering method outlined in ProtBFN (Atkinson et al., 2025) with 32 particles. More details on IgCraft’s sampling methodology can be found in Appendix A.
@@ 2.3 Datasets and training regime
Training IgCraft consists of three stages. First, each chain’s transformer stack (Figure 1, blue) is pre-trained on unpaired sequences. Then, the model is fine-tuned on paired sequences, for which the pre-trained weights for both transformer stacks are loaded into the network and the interaction blocks (Figure 1, green) are randomly initialized. However, we initialize the bias term in the output gate of each interaction block to a value of , negating its contribution to the token embeddings at the start of fine-tuning and effectively initializing the model as two unpaired sequence models that do not communicate. All weights are updated during this stage of fine-tuning, including the weights initialized from pre-training. Finally, the model is fine-tuned using paired antibody structures as conditioning information for the model’s sequence predictions, initializing the output gate of the structure encoder to . During structure-fine-tuning, the framework regions of input structures are masked stochastically in 50% of training examples to train the model to perform CDR-conditional framework generation. In this stage, only the weights in the structure encoder are updated to ensure that the main trunk retains its capabilities as a pure sequence generative model. This approach is conceptually similar to the approach proposed by Zheng et al. (2023) for performing inverse folding by augmenting a protein language model with a lightweight structural adapter.
@@ 3.3 Inverse folding and structure-guided sequence design
To assess the model’s ability to conditionally generate antibody sequences in the presence of structural information, we sampled a single sequence for each of the 98 test structures using IgCraft with a temperature of 0.1, providing the antibody and target backbone structures as input to the model’s structure encoder but exposing no sequence information from the antibody (target sequence identities are used as input features). For comparison, we also performed inverse folding (also using temperature 0.1) using ProteinMPNN (Dauparas et al., 2022), AbMPNN (Dreyer et al., 2023), and Antifold (Høie et al., 2024). While the target was cropped to 128 residues when sampling with IgCraft, the full target chains were provided as input to all other inverse folding methods. To measure performance, we calculate the amino acid recovery for each variable domain region and estimate two properties related to the developability potential of the generated sequences: humanness and solubility. Humanness scores for the generated sequences are obtained using AbNatiV (Ramon et al., 2024) while solubility is scored with CamSol (Sormanni et al., 2015; Rosace et al., 2023), both of which have been supported by experimental validation. Results are shown in Table 3.
@@ 3.4 CDR grafting and humanisation
Scaffolding (”grafting”) of CDRs from mouse or other non-human antibodies into human antibody framework regions is a key task in antibody engineering (Jones et al., 1986; Kim & Hong, 2012), since non-human antibodies are often easier to obtain but can induce pathological anti-drug immune responses when administered as therapeutics (Khazaeli et al., 1994). Traditional CDR grafting workflows typically rely on sequence homology searches to select an existing human framework similar to the input antibody (Kim & Hong, 2012). However, these approaches are inherently limited by the availability of similar framework sequences and often substantially decrease binding affinity by altering key interactions between the CDR and framework regions (Pavlinkova et al., 2001). HuDiff was recently proposed as an ML-driven solution for this task (Ma et al., 2024), which uses a discrete diffusion model to conditionally generate framework sequences given a set of input CDR sequences, and is fine-tuned specifically on mouse antibodies to sample mutations that improve the AbNatiV humanness score (Ramon et al., 2024) of generated sequences. The authors include impressive experimental evidence demonstrating that the binding properties of an existing high-affinity mouse antibody (<1nM kD) were largely preserved after humanisation. However, HuDiff provides no mechanism for integrating structural information for the input CDRs.
@@
Given IgCraft’s strong performance in generating human framework sequences and its capacity to condition on structural inputs, we sought to evaluate the model’s ability to generate human framework sequences for scaffolding mouse CDRs. We performed conditional sampling with IgCraft using 27 mouse antibodies from SAbDab as input and generating framework sequences for both chains, providing the CDR structures as conditioning information as well as the CDR sequences 2 residues on each side. For each set of mouse CDRs, we also sampled a single paired sequence from HuDiff (Ma et al., 2024) as a benchmark. The effectiveness of CDR grafting is measured in two ways: first, the extent to which the humanness of the antibody is improved after grafting (compared to the parental antibody), and second, the extent to which the binding properties of the antibody are preserved. We evaluated humanness using AbNatiV (Ramon et al., 2024), an autoencoder-based deep learning method, and OASis (Prihoda et al., 2022), an approach that estimates humanness using 9-mer peptide frequences in OAS. To test how well both grafting methods maintained the structure and binding properties of the wild-type CDRs, we applied AlphaFold3 (Abramson et al., 2024) to fold each generated sequence (and its target, if applicable) with 10 seeds, as well as the 27 original mouse sequences for comparison. For each structure prediction we measure the RMSD of the H-CDR3 loop (superimposing only on the framework region), and for bound antibodies, determine whether the antibody was docked to the correct epitope on the target protein using the widely-applied DockQ score threshold of 0.23 (Mirabello & Wallner, 2024; Abramson et al., 2024). Results are shown in Table 4. We also include an ablation study in Appendix A in which CDR grafting was performed using IgCraft without structural information, demonstrating that providing CDR structures as input significantly improves the model’s ability to propose framework sequences that preserve the structural features of the input CDRs.
@@ A.1.3 Conditional sampling
For conditional sampling tasks with IgCraft we implement the sequential monte carlo (SMC) framework from ProtBFN (Trippe et al., 2022; Atkinson et al., 2025). In general, SMC methods attempt to approximate a posterior distribution over the data given some conditioning information :
In sequential importance resampling, at each step of the sampling trajectory, proposal samples are first drawn from the unconditional model and then re-sampled with replacement under some appropriately normalized likelihood function . The trained BFN - via the SDE in (1) - provides access to the unconditional distribution over token logits . The problem of sequence inpainting considers the task of conditioning generation on a subset of sequence positions with corresponding tokens . Denoting samples of the proposal distribution as , ProtBFN (Atkinson et al., 2025) uses the following form for the conditional likelihood:
@@ A.1.4 Gated residual connections
A key element of IgCraft’s network architecture is the use of gated residual connections (Srivastava et al., 2015; Dhayalkar, 2024) at specific positions between layers that are updated at different stages of the training process (Figure 1). A standard sigmoid-gated linear unit (GLU) (Dauphin et al., 2016) is defined as follows: